OpenStudy (anonymous):

1. Sally said that when you add or subtract positive or negative decimals, you can change the order of the decimals and still get the same answer. Use these decimals to answer the questions: –6.2, 5.71, –3.96, 2.84 (a) Is Sally’s statement always true, sometimes true, or always false? Write two pairs of equations using addition and two pairs of equations using subtraction to support your answer. (b) Write a rule about changing the order of the numbers in adding and in subtracting positive and negative decimals. (c) Is Sally’s statement true for addition and subtraction of three pos

OpenStudy (anonymous):

i will medal and fan

OpenStudy (anonymous):

I think that a is sometimes true

OpenStudy (anonymous):

but im not sure

OpenStudy (anonymous):

I think it's always true

OpenStudy (anonymous):

thx and the rest of c is: (c) Is Sally’s statement true for addition and subtraction of three positive or negative decimals? Write two pairs of equations (one pair for addition, and one pair for subtraction) that support your answer.

OpenStudy (anonymous):

-7+(-3)=-10 -3+(-7)=-10 -7-(-3)=-4 -3-(-7)=4 Its sometimes true

OpenStudy (anonymous):

sooooooooo

OpenStudy (anonymous):

@DancerAtHeart111 am I right???

OpenStudy (anonymous):

yes thanks so much!!!